# MLRC-Bench: Can Language Agents Solve Machine Learning Research Challenges?

### v1

@@ Acknowledgements
This work is supported by LG AI Research.
Impact Statement
The ability of AI agents to perform high-quality ML research could accelerate breakthroughs in critical domains such as healthcare, climate modeling, and AI safety. However, agents capable of autonomously generating novel methods at scale may also amplify risks if their outputs outpace human understanding or oversight. While current agents underperform human researchers, MLRC-Bench highlights the need for ongoing monitoring of their capabilities to ensure alignment with ethical standards and societal goals.
To mitigate risks, MLRC-Bench emphasizes tamper-proof evaluation protocols and objective performance metrics, reducing reliance on subjective judgments that may overestimate innovation. By releasing this benchmark, we aim to enhance transparency and encourage the development of safer, more reliable AI research agents. We caution against deploying such systems without robust safeguards and urge the community to prioritize evaluations that balance innovation with accountability.
@@ Appendix A Detailed Description of Research Competitions
A.2 Backdoor Trigger Recovery (Xiang et al., 2024) [Link]
@@
A.3 Temporal Action Localisation (Heyward et al., 2024) [Link]
@@
A.4 Rainfall Prediction (Gruca et al., 2022) [Link]
@@
A.5 Machine Unlearning (Triantafillou et al., 2024) [Link]
@@
A.6 Next Product Recommendation (Jin et al., 2023) [Link]
@@
A.7 Cross-Domain Meta Learning (Carrión-Ojeda et al., 2022) [Link]
@@
Appendix B Additional Results
This section presents additional results that complement the findings reported in the main paper.
@@ Appendix B Additional Results
  • Figure 9 shows inference-time scaling results broken down by task, complementing the aggregate results in Figure 2 (Section 4.3).
  • Figures 10 and 11 provide a task-level analysis of the implementation process, extending the results in Figure 5 (Section 4.5).
@@
Human Idea + MLAB (gpt-4o) 1.5 -0.7 -14.9 8.9 26.1 4.2 54.5 11.4
CoI-Agent Idea (o1) + MLAB (gpt-4o) 1.0 -0.7 -14.9 0.6 83.6 7.3 24.9 14.5
@@
Appendix C Prompts for LLM-as-a-Judge
@@
Appendix D Agent Trace Analysis
@@
D.1 Error Type Categorization
@@
  • Non-execute Steps: Steps where the action ”Execute Script” was not invoked.
  • Execute Steps: Steps where the action ”Execute Script” was invoked.
@@
Figure 12 shows that Gemini-exp-1206 successfully completes a considerable number of steps without errors, yet its performance varies noticeably across tasks. In particular, while Meta Learning displays relatively few issues, Rainfall Prediction exhibits a higher frequency of ”hallucination” based errors such as incorrect argument handling, non-existent file references, and invalid parameter choices. This discrepancy indicates that certain tasks present greater challenges for the model, likely due to more complex or less familiar contexts.
Within the Execute Steps, the most frequent error types are import, value, and type errors, reflecting a tendency to reference nonexistent modules, pass invalid parameters, or supply arguments of the wrong data type. On the Non-execute Steps side, incorrect arguments remain a recurring challenge, showing another case where the agent seems to be ”hallucinating” the argument names.
@@
D.2 Error Response Distribution
@@
From Figure 13, we observe notable variations in error-handling effectiveness across tasks. Specifically, LLM Merging demonstrates the highest proportion of fully resolved (Fixed) errors, indicating more effective resolution strategies, whereas Rainfall Prediction, Backdoor Trigger Recovery, Machine Unlearning, and Perception Temporal Action Localization predominantly exhibit errors that are only partially addressed (Attempted). Meanwhile, Meta Learning has the largest share of errors categorized as Did not attempt. These distinctions highlight task-specific differences in error management
We also observe a consistently high percentage of errors categorized as Attempted across nearly all tasks, indicating that the agent often struggles to fully resolve errors. This broadly suggests challenges in the agent’s comprehension or planning capabilities when addressing complex errors, potentially pointing to difficulties in fully interpreting the underlying problem or effectively formulating corrective actions. Additionally, the notable variability in fully resolved (Fixed) and unaddressed (Did not attempt) errors across tasks implies that certain tasks inherently pose greater cognitive complexity or ambiguity, further exacerbating the agent’s difficulty in error resolution. The prompts used for this annotation are shown in Appendix G.
@@
D.3 Error Solve Rate
To further expand on this analysis of errors, we also show which error types are more effectively resolved and highlight their associated complexity during task execution. Errors which were categorised as Fixed are treated as solved while errors belonging to the other two categories are treated as unsolved. Using the prompt in Appendix G, we also had GPT-4o-mini return the step at which the error was fixed for those that were categorised as fixed.
@@ D.3 Error Solve Rate
D.4 Per-Step Action Distribution
@@
D.5 Per-Step Stage Distribution
@@
Toward the latter steps, a subset of runs proceeds to Experimental Refinement, engaging in repeated re-runs, parameter tuning, and exploring alternative strategies to optimize performance. However, in many cases, the agent transitions relatively quickly to Final Evaluation & Submission. This early move towards final submission implies potential underuse of iterative enhancement cycles, indicating an area for improvement in the agent’s approach. The prompts used for this annotation are shown in Appendix F.
D.6 Stage Timelines
@@ D.7 Capability Level
Appendix E Agent Code Samples
Here we show the two examples of solutions generated by LLM agents mentioned in Section 4.6. Please see all the agent code in our benchmark codebase [^7] https://github.com/yunx-z/MLRC-Bench.
@@ 45 The unlearned model
Appendix F Prompt for Stage Annotation
@@ 7. Final Evaluation & Submission:
IMPORTANT: The original trace has {original_step_count} steps. Your response MUST contain exactly {original_step_count} keys, numbered from ”1” to ”{original_step_count}”.
@@
Appendix G Prompt for Error Response Annotation

### v2

@@ Acknowledgements
This work is supported by LG AI Research. We thank Xinnuo Li and Aryan Sharma for their help with the survey and collection of research competitions. We also thank LAUNCH lab members for their useful feedback.
@@ A.1 LLM Merging [47] [Link]
A.2 Backdoor Trigger Recovery [51] [Link]
@@
A.3 Temporal Action Localisation [18] [Link]
@@
A.4 Rainfall Prediction [17] [Link]
@@
A.5 Machine Unlearning [48] [Link]
@@
A.6 Next Product Recommendation [26] [Link]
@@
A.7 Cross-Domain Meta Learning [6] [Link]
@@
Appendix B Additional Discussion on Related Work
Table 3: Comparison between MLRC-Bench and existing work on automated scientific discovery in machine learning with LLM agents. $\sim$” means that some but not all of the tasks in that benchmark require the indicated capability. “Compute Constraints” indicates whether the solution code must adhere to specified runtime and GPU memory limitations.
Problem Identification Method Proposal Experiment Design Code Implementation Evaluation Method Evaluation Object Compute Constraints Continual Updates
AI Scientist [36] LLM & Human Judge Paper
Can LLMs Generate Novel Research Ideas? [44] Human Judge Idea Proposal
DiscoPOP [35] Performance -Based Function-Level Code
MLAgentBench [20] $\sim$ Performance -Based File-Level Code
MLE-Bench [7] $\sim$ Performance -Based File-Level Code
MLGym-Bench [39] $\sim$ Performance -Based File-Level Code
RE-Bench [50] Performance -Based Repository -Level Code
MLRC-Bench (Ours) Performance -Based Repository -Level Code
Scientific discovery in machine learning typically includes four main stages: Problem Identification, where gaps in existing methods are recognized; Method Proposal, which introduces a new approach to address the issue; Experiment Design, involving the selection of datasets, baselines, and metrics for evaluation; and Code Implementation, where the method is realized through executable code. While prior work [36, 44] covers Problem Identification and Experiment Design, evaluation could be subjective based on idea proposal or final paper. Instead, MLRC-Bench focuses on the critical stages of proposing and implementing novel methods, enabling objective performance assessment. The differences between our benchmark and existing work on automating ML research workflow with LLM agents are presented in Table 3.
Appendix C Additional Analysis
C.1 Case Study
We present two case studies below to illustrate failed solutions implemented by LLM agents. Please see Appendix G for the concrete code implemented by AI agents.
Machine Unlearning Challenge: The goal is to develop efficient algorithms that enable a model to “forget" specific training data, such that the resulting model closely resembles one that was never trained on that data in the first place. The MLAB agent (claude-3-5-sonnet-v2) proposed a Gradient Ascent Unlearning Algorithm, a two-phase approach combining gradient ascent for forgetting and fine-tuning for retaining knowledge. Specifically, the algorithm first performs gradient ascent on the forget set to maximize loss (achieving unlearning) and then fine-tunes the model on the retain set to restore the desired knowledge. While this approach sounds promising in theory, it scored significantly lower than the baseline. We hypothesize that by separating the gradient ascent on the forget set and the fine-tuning on the retain set into two distinct phases, the model may not effectively balance these two conflicting objectives. In contrast, a joint optimization approach—where both objectives are optimized at each gradient update—might better balance the processes of “forgetting” and “retaining” knowledge.
Figure 6: We perform a cost-effectiveness analysis of various setups. On the x-axis, we plot API cost, where lower is better, and on the y-axis, we show reletive improvement to human (Section 3.3), where higher is better.
C.2 Cost-Effectiveness Analysis
In Figure 6, we analyze the agents’ success rates in a cost-controlled setting, motivated by recent work [28] emphasizing the importance of jointly optimizing both performance and cost in agent design. [^9] We exclude the gemini-exp-1206 model from this figure because it was experimental and its pricing was unavailable at the time of writing. Llama 3.1 405b Instruct [^10] We estimate the API cost for Llama models based on Amazon Bedrock service pricing. offers the most favorable trade-off, achieving higher success rates than GPT-4o and Claude 3.5 Sonnet at a significantly lower cost.
Appendix D Additional Results
This section presents additional results that complement the findings reported in the main paper and appendix.
@@ Appendix D Additional Results
  • Figure 8 shows inference-time scaling results broken down by task, complementing the aggregate results in Figure 5 (Appendix 4.5).
  • Figures 9 and 10 provide a task-level analysis of the implementation process, extending the results in Figure 2 (Appendix 4.3).
@@
w/ Human Idea 1.5 -0.7 -14.9 8.9 26.1 4.2 54.5 11.4
w/ CoI-Agent Idea (o1) 1.0 -0.7 -14.9 0.6 83.6 7.3 24.9 14.5
@@
Appendix E Prompts for LLM-as-a-Judge
@@
Appendix F Agent Trace Analysis
@@
F.1 Error Type Categorization
@@
  • Non-execute Steps: Steps where the action "Execute Script" was not invoked.
  • Execute Steps: Steps where the action "Execute Script" was invoked.
@@
Figure 11 shows that Gemini-exp-1206 successfully completes a considerable number of steps without errors, yet its performance varies noticeably across tasks. In particular, while Meta Learning displays relatively few issues, Rainfall Prediction exhibits a higher frequency of "hallucination" based errors such as incorrect argument handling, non-existent file references, and invalid parameter choices. This discrepancy indicates that certain tasks present greater challenges for the model, likely due to more complex or less familiar contexts.
Within the Execute Steps, the most frequent error types are import, value, and type errors, reflecting a tendency to reference nonexistent modules, pass invalid parameters, or supply arguments of the wrong data type. On the Non-execute Steps side, incorrect arguments remain a recurring challenge, showing another case where the agent seems to be "hallucinating" the argument names.
@@
F.2 Error Response Distribution
@@
From Figure 12, we observe notable variations in error-handling effectiveness across tasks. Specifically, LLM Merging demonstrates the highest proportion of fully resolved (Fixed) errors, indicating more effective resolution strategies, whereas Rainfall Prediction, Backdoor Trigger Recovery, Machine Unlearning, and Perception Temporal Action Localization predominantly exhibit errors that are only partially addressed (Attempted). Meanwhile, Meta Learning has the largest share of errors categorized as Did not attempt. These distinctions highlight task-specific differences in error management.
We also observe a consistently high percentage of errors categorized as Attempted across nearly all tasks, indicating that the agent often struggles to fully resolve errors. This broadly suggests challenges in the agent’s comprehension or planning capabilities when addressing complex errors, potentially pointing to difficulties in fully interpreting the underlying problem or effectively formulating corrective actions. Additionally, the notable variability in fully resolved (Fixed) and unaddressed (Did not attempt) errors across tasks implies that certain tasks inherently pose greater cognitive complexity or ambiguity, further exacerbating the agent’s difficulty in error resolution. The prompts used for this annotation are shown in Appendix I.
@@
F.3 Error Solve Rate
To further expand on this analysis of errors, we also show which error types are more effectively resolved and highlight their associated complexity during task execution. Errors which were categorised as Fixed are treated as solved while errors belonging to the other two categories are treated as unsolved. Using the prompt in Appendix I, we also had GPT-4o-mini return the step at which the error was fixed for those that were categorised as fixed.
@@ F.3 Error Solve Rate
F.4 Per-Step Action Distribution
@@
F.5 Per-Step Stage Distribution
@@
Toward the latter steps, a subset of runs proceeds to Experimental Refinement, engaging in repeated re-runs, parameter tuning, and exploring alternative strategies to optimize performance. However, in many cases, the agent transitions relatively quickly to Final Evaluation & Submission. This early move towards final submission implies potential underuse of iterative enhancement cycles, indicating an area for improvement in the agent’s approach. The prompts used for this annotation are shown in Appendix H.
F.6 Stage Timelines
@@ F.7 Capability Level
Appendix G Agent Code Samples
Here we show the two examples of solutions generated by LLM agents mentioned in Section C.1.
@@ 45 The unlearned model
Appendix H Prompt for Stage Annotation
@@ 7. Final Evaluation & Submission:
IMPORTANT: The original trace has {original_step_count} steps. Your response MUST contain exactly {original_step_count} keys, numbered from "1" to "{original_step_count}".
@@
Appendix I Prompt for Error Response Annotation
@@
Appendix J Impact Statement
The ability of AI agents to perform high-quality ML research could accelerate breakthroughs in critical domains such as healthcare, climate modeling, and AI safety. However, agents capable of autonomously generating novel methods at scale may also amplify risks if their outputs outpace human understanding or oversight. While current agents underperform human researchers, MLRC-Bench highlights the need for ongoing monitoring of their capabilities to ensure alignment with ethical standards and societal goals. By releasing this benchmark, we aim to enhance transparency and encourage the development of safer, more reliable AI research agents. We caution against deploying such systems without robust safeguards and urge the community to prioritize evaluations that balance innovation with accountability.